OpenStudy (anonymous):

Could spomeone please take the time to help me understand this?? Write the point-slope form of the equation passing through (5, -1) with a slope of 6.

OpenStudy (anonymous):

these are the answers : y-1=6(x+5) y+5=6(x-1) y+1=6(x-5) y-5=6(x+1)

OpenStudy (anonymous):

point slope form y-y1=m(x-x1) m=slope

OpenStudy (anonymous):

y-(-1)=6(x-5) y+1=6(x-5)

OpenStudy (anonymous):

there is only one set of numbers how can i find m?

OpenStudy (anonymous):

slope is given in question see it

OpenStudy (anonymous):

oh so m=6

OpenStudy (anonymous):

any more question about this

OpenStudy (kymber):